# Lifelong Learning of Video Diffusion Models From a Single Video Stream

### v2

@@ Appendix E Training Time Performance Metrics
(d) Lifelong PLAICraft dataset.
Figure 8: Test stream performance metrics for model checkpoints at different training iterations. The plots show the improvement in model quality as online training progresses. All models generally improve the longer they are trained.
Appendix F Train Stream Performance Breakdown
@@
(d) Lifelong PLAICraft dataset.
Figure 9: Fully trained models’ performance metrics on video frames from different parts of the training stream. The plots show whether the final model performs better or worse on future frame prediction for frames earlier or later in the training stream. Each point of the plots is calculated using 1,000 consecutive video frame subsequences that succeed the corresponding train stream frame index.
Appendix G Additional Qualitative Results
G.1 Lifelong Bouncing Balls
G.1.1 Train Stream Examples
Figure 10: Lifelong Bouncing Balls (O) trajectories from a video early in the training stream.
Figure 11: Lifelong Bouncing Balls (O) trajectories from a video late in the training stream.
Figure 12: Lifelong Bouncing Balls (C) trajectories from a video early in the training stream.
Figure 13: Lifelong Bouncing Balls (C) trajectories from a video late in the training stream.
G.1.2 Test Stream Examples
Figure 14: Lifelong Bouncing Balls (O) trajectories from a video early in the test stream.
Figure 15: Lifelong Bouncing Balls (O) trajectories from a video late in the test stream.
Figure 16: Lifelong Bouncing Balls (C) trajectories from a video early in the test stream.
Figure 17: Lifelong Bouncing Balls (C) trajectories from a video late in the test stream.
G.2 Lifelong 3D Maze
G.2.1 Train Stream Examples
Figure 18: Lifelong 3D Maze crystal sequence from the training stream.
Figure 19: Lifelong 3D Maze painting scenario from the training stream.
G.2.2 Test Stream Examples
Figure 20: Lifelong 3D Maze crystal sequence from the test stream.
Figure 21: Lifelong 3D Maze painting scenario from the test stream.
G.3 Lifelong PLAICraft
G.3.1 Train Stream Examples
Figure 22: Crafting table scenario.
Figure 23: Chest scenario.
Figure 24: Looking at another player in front of the village scenario.
Figure 25: Walking forward on a plain scenario.
Figure 26: Walking forward by a mountain on the side scenario.
Figure 27: Sleeping scenario.
Figure 28: Walking forward in The End biome scenario.
G.3.2 Test Stream Examples
Figure 29: Crafting table scenario.
Figure 30: Chest scenario.
Figure 31: Walking forward on a plain area near some tall buildings scenario.
Figure 32: Swimming in an ocean with the f3 information menu on scenario.

### v3

@@ Appendix F Optimization Time Performance Metrics
(d) Lifelong Drive dataset.
(e) Lifelong PLAICraft dataset.
Figure 5: Test stream performance metrics for the U-Net model checkpoints at different training iterations (refer to Appendix D for baseline details). The plots show the improvement in model quality as online training progresses. All models generally improve the longer they are trained.
Appendix G Train Stream Performance Breakdown
@@
(d) Lifelong Drive dataset.
(e) Lifelong PLAICraft dataset.
Figure 6: Fully trained U-Net models’ performance metrics on video frames from different parts of the training stream (refer to Appendix D for baseline details). The plots show whether the final model performs better or worse on future frame prediction for frames earlier or later in the training stream. Each point of the plots is calculated using 1,000 consecutive video frame subsequences that succeed the corresponding train stream frame index.